# Wikontic: Constructing Wikidata-Aligned, Ontology-Aware Knowledge Graphs with Large Language Models

### v1

@@ 3.2 Graph Quality Analysis
Method $\mathcal{E}|$ $\mathcal{R}|$
Avg. $e$
degree
Unique
$e$ per $r$
$r$ diversity
per $2\times e$
@@
w/o ontology (2)
232.4 106.7 4.4 2.6 1.06
w/o ontology (2) and
    normalization (3)
273.0 140.9 4.2 2.3 1.09
w/o filtered triplets
239.9 99.5 4.3 2.6 1.0
@@
A KG can be formally represented as $\mathcal{G}=(\mathcal{T},\mathcal{E},\mathcal{R})$, where $\mathcal{E}$ is the set of entities $e$, $\mathcal{R}$ is the set of relations $r$ and $\mathcal{T}$ is the set of triplets: $\mathcal{T}\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}$. For efficient knowledge storage and retrieval, the KG should satisfy the size, density and diversity desiderata, which can be directly evaluated using graph statistics (Table 2).
@@
High KG connectivity, represented by the average entity degree, ensures efficient retrieval, especially with limited search depth. Entity normalization (Stage 3) is a key to building KGs with the highest density among the compared methods. The relation diversity, or the number of unique relations per two entities, captures a variety of information stored in the KG. Here, Wikontic versions with relaxed ontology constraint remain the most diverse.
@@
3.3 Answer Coverage
To measure the overall coverage of various KG construction methods, we estimate whether the answer to the question is present in the KG as an entity and does the path from question to answer exist in the graph. Due to differences in pipelines, we cast all entity and relation names to lowercase and remove punctuation to standardize their format. To account for possible differences in entity naming, we consider two entities matching if one is a substring of another.
Table 3 presents coverage and size metrics for KGs built by our pipeline, AriGraph and HippoRAG on samples from the MuSiQue dataset. Since baseline KGs are available only for gpt4o-mini model and 80 common test samples, we use the same configuration to ensure a fair comparison. We estimated the standard deviation using bootstrapping. "Contains Answer" represents the percentage of cases when the answer entity is present in the full KG, or 5- and 10- hop neighborhood of entities present in MuSiQue questions. We ablate Wikontic by removing one or multiple pipeline steps at a time. "Ontology Entailment" is the percentage of triplets, where subject, object and relation match the ontology.
Method Contains Answer (%)
Ontology
Entailment (%)
@@ 3.3 Answer Coverage
Table 3: Knowledge coverage of graphs built using different extraction methods on the MuSiQue dataset with mini-sized models. (Left) Percentage of cases where the correct answer to a question appears in the full constructed graph or within the 5-, and 10-hop neighborhoods of the question nodes. Wikontic pipelines provide the best answer coverage while maintaining high ontology agreement. (Right) Average percentage of triplets in each sample that are entailed by the Wikidata ontology.
@@
ReadAgent lee2024human, gpt4 35.0 45.1 48.9 62.0
GraphReader li2024graphreader, gpt4 38.0 47.4 55.0 70.0
GraphRAG edge2024local, gpt4o-mini 40.0 53.5 58.7 63.3
AriGraph anokhin2024arigraph, gpt4o-mini 36.5 47.9 60.0 68.0
AriGraph anokhin2024arigraph, gpt4 45.0 57.0 68.0 74.7
HOLMES panda2024holmes, gpt4 48.0 58.0 66.0 78.0
@@
HippoRAG v2 gutierrez2025rag, Llama 3.3 37.2 48.6 62.7 75.5
@@
Both Wikontic and HippoRAG achieve over 96% answer coverage, surpassing AriGraph’s 79.9%. Notably, only 3.5% of triples in Wikontic are ontology-misaligned, confirming that the generated knowledge is largely schema-consistent, only a small portion of produced triplets requires correction to fully satisfy ontology constraints. Without ontology constraints, Wikontic reaches the highest coverage (97.5%), avoiding mismatches between Wikidata and MuSiQue entities and making it particularly effective for open-domain QA. When ontology consistency is required, the filtered Wikontic variant attains 100% ontology entailment with a competitive 93.8% coverage.
These findings suggest that Wikontic variants provide the best solution both with and without the ontology. Ontology and filtering help achieve the most standardized graph and strong answer coverage by thorough triplet refinement and deduplication strategies, Wikontic achieves the best coverage of information essential for question answering, while maintaining the highest connectivity levels, crucial for enabling graph search.
@@ 3.4 Computational Efficiency
A key indicator of computational cost is the number of completion tokens, which are typically around 3-5 times more expensive [^4] https://claude.com/platform/api/ [^5] https://openai.com/api/pricing/ and computationally intensive than input tokens zhou2024survey. Under this metric, Wikontic is more efficient, producing KGs with roughly three times fewer output tokens than AriGraph (881 vs 2,500) and about twenty times fewer than GraphRAG (881 vs 20,000). This shows that Wikontic achieves comparable KG construction quality while using significantly fewer output tokens.
@@ 3.5 Performance on QA Tasks
Despite this constraint, using our triplet-only contexts for gpt-4.1, it achieves strong results, reaching 64.5 EM and 76.0 F1 on HotpotQA and 46.8 EM and 59.8 F1 on MuSiQue. These scores surpass several retrieval-based methods, including ReadAgent and GraphReader, and are comparable to more resource-intensive systems such as AriGraph and HOLMES that rely on more rich textual context. This demonstrates that complete and well-structured symbolic representations of KGs can serve as a sufficient and reliable information source for multi-hop reasoning.
@@ 4 Conclusions
We introduced Wikontic, a fully automated pipeline that uses LLMs to construct KGs from unstructured text. The pipeline produces compact, internally consistent graphs by aligning extracted triplets with the Wikidata ontology and deduplicating entities and relations. While using KGs as the sole knowledge source for multi-hop question answering, Wikontic achieves competitive performance with retrieval-augmented and KG-based baselines that still rely on source texts, demonstrating that ontology-guided KG construction is a viable alternative to passage-level retrieval. On MuSiQue, Wikontic includes 38–45 more unique entities than HippoRAG and AriGraph and contains the correct answer entity in 97.5% of cases. Within a 10-hop subgraph, it maintains 70% answer coverage and denser local connectivity. For QA, it attains 64.5 EM / 76.0 F1 on HotpotQA and 46.8 EM / 59.8 F1 on MuSiQue, outperforming text-reliant systems (ReadAgent, GraphReader) and approaching larger text-dependent methods (AriGraph, HOLMES). Moreover, Wikontic achieves state-of-the-art results on the MINE-1 benchmark, achieving 84–86% information-retention scores and surpassing GraphRAG and KGGen. These findings indicate that the proposed pipeline preserves a substantial amount of factual information from source texts.
@@
Our findings show that ontology-aware KG construction enables scalable, interpretable, and verifiable transformation of unstructured text into structured knowledge, bridging together symbolic reasoning and generative language modeling.
@@
Our experiments are restricted to proprietary OpenAI models (GPT-4.1, GPT-4.1-mini, GPT-4o-mini) and the open-source Llama 3.3-70B. Token efficiency is measured as model-generated tokens during answer production. This metric reflects provider billing but not end-to-end latency or throughput. Every stage of the pipeline currently uses LLMs with instructions and in-context examples prompting. Because the pipeline now yields its own annotated data, several stages could be replaced in future work by smaller, task-specific models fine-tuned on this data, thereby improving efficiency and lowering computational cost.
@@ A.1 Reproducibility
We release: (1) prompts and source code for each pipeline stage; (2) scripts to build Wikidata-derived ontology used for ontology-aware stages; (3) the KG-only multi-hop QA component. We also plan to release all KGs that were build with Wikontic pipeline. The source code is included as Supplementary Material with this submission.
@@ A.2 Dataset Statistics
The test splits of HotpotQA and MuSiQue consist of 1000 samples each. We provide the QA evaluation results for Wikontic and HippoRAG for the whole evaluation set and for AriGraph we report the openly available results for 200 test samples. To compare the statistics of KGs we use 80 MuSiQue samples that are commonly available for all compared methods.
@@ A.4 Prompts for triplet extraction
Here we provide excerpts of prompts that were used in our KG construction pipeline. A.4 was used for candidate triplet extraction. Subsequently, A.4 was used to refine entity types for both subject and object entities. A.4 was used for choosing relation among ones that can legally connect chosen entity types by Wikidata constraints. Finally, A.5 was used to refine surface forms of subject and object. All prompts, instructions, in-context examples are available with the code.
@@ A.5 Prompts for question answering
Here we provide excerpts of prompts that were used for KG grounded question answering. A.5 was used to extract entities relevant for the question. Then, with A.5 the LLM was instructed to choose relevant entities among entities similar to the extracted ones in the KG. A.5 was used to decompose the question on single-hop subquestion conditioned on extracted entities or previously answered subquestion. A.5 was used to check if a question is answered by a sequence of subquestions and corresponding answers. All prompts, instructions, in-context examples are available with the code.

### v2

@@ 3.2 Graph Quality Analysis
A KG can be formally represented as $\mathcal{G}=(\mathcal{T},\mathcal{E},\mathcal{R})$, where $\mathcal{E}$ is the set of entities $e$, $\mathcal{R}$ is the set of relations $r$ and $\mathcal{T}$ is the set of triplets: $\mathcal{T}\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}$. For efficient knowledge storage and retrieval, the KG should satisfy the size, density, and diversity desiderata, which can be directly evaluated using graph statistics (Table 2).
Method $\mathcal{E}|$ $\mathcal{R}|$ Avg. $e$ degree Unique $e$ per $r$ $r$ diversity per $2\times e$
@@
w/o ontology (2) 232.4 106.7 4.4 2.6 1.06
w/o ontology (2) and     normalization (3) 273.0 140.9 4.2 2.3 1.09
w/o ontology-misaligned     triplets 239.9 99.5 4.3 2.6 1.0
@@
High KG connectivity, represented by the average entity degree, ensures efficient retrieval, especially with limited search depth. Entity normalization (Stage 3) is a key to building KGs with the highest density among the compared methods. The relation diversity, or the number of unique relations per two entities, captures a variety of information stored in the KG. Here, Wikontic versions with relaxed ontology constraints remain the most diverse.
@@
Method Contains Answer (%) Ontology Entailment (%)
@@
Table 3: Knowledge coverage of graphs built using different extraction methods on the MuSiQue dataset with mini-sized models. (Left) Percentage of cases where the correct answer to a question appears in the full constructed graph or within the 5-and 10-hop neighborhoods of the question nodes. Wikontic pipelines provide the best answer coverage while maintaining high ontology agreement. (Right) Average percentage of triplets in each sample that are entailed by the Wikidata ontology.
@@
ReadAgent Lee et al. (2024), gpt4 35.0 45.1 48.9 62.0
GraphReader Li et al. (2024), gpt4 38.0 47.4 55.0 70.0
GraphRAG Edge et al. (2024), gpt4o-mini 40.0 53.5 58.7 63.3
AriGraph Anokhin et al. (2025), gpt4o-mini 36.5 47.9 60.0 68.0
AriGraph Anokhin et al. (2025), gpt4 45.0 57.0 68.0 74.7
HOLMES Panda et al. (2024), gpt4 48.0 58.0 66.0 78.0
@@
HippoRAG v2 Gutiérrez et al. (2025), Llama 3.3 37.2 48.6 62.7 75.5
@@
3.3 Answer Coverage
To measure the overall coverage of various KG construction methods, we estimate whether the answer to the question is present in the KG as an entity and whether the path from the question to the answer exists in the graph. Due to differences in pipelines, we cast all entity and relation names to lowercase and remove punctuation to standardize their format. To account for possible differences in entity naming, we consider two entities matching if one is a substring of the other.
Table 3 presents coverage and size metrics for KGs built by our pipeline, AriGraph, and HippoRAG on the MuSiQue dataset. Since baseline KGs are available only for the gpt4o-mini model and 80 common test samples, we use the same configuration to ensure a fair comparison. We estimated the standard deviation using bootstrapping. "Contains Answer" represents the percentage of cases when the answer entity is present in the neighborhood of entities from the MuSiQue question. We ablate Wikontic by removing one or multiple pipeline steps at a time. "Ontology Entailment" is the percentage of triplets where subject, object, and relation match the ontology.
Both Wikontic and HippoRAG achieve over 96% answer coverage, surpassing AriGraph’s 79.9%. Notably, only 3.5% of triples in Wikontic are ontology-misaligned, confirming that the generated knowledge is largely schema-consistent; only a small portion of produced triplets requires correction to fully satisfy ontology constraints. Without ontology constraints, Wikontic reaches the highest answer coverage (97.5%), avoiding mismatches between Wikidata and MuSiQue entities and making it particularly effective for open-domain QA. When ontology consistency is required, the Wikontic variant that excludes ontology-misaligned triplets (100% ontology entailment) attains a competitive answer coverage of 93.8%.
These findings suggest that Wikontic variants provide the best solution both with and without the ontology. Ontology and misaligned triplets exclusion help to achieve the most standardized graph and strong answer coverage by thorough triplet refinement and deduplication strategies. Wikontic achieves the best coverage of information essential for question answering, while maintaining the highest connectivity levels, crucial for enabling graph search.
@@ 3.4 Computational Efficiency
A key indicator of computational cost is the number of completion tokens, which are typically around 3-5 times more expensive [^4] https://claude.com/platform/api/, [^5] https://openai.com/api/pricing/ and computationally intensive than input tokens Zhou et al. (2024). Under this metric, Wikontic is more efficient, producing KGs with roughly three times fewer output tokens than AriGraph (881 vs 2,500) and about twenty times fewer than GraphRAG (881 vs 20,000). This shows that Wikontic achieves comparable KG construction quality while using significantly fewer output tokens.
@@ 3.5 Performance on QA Tasks
Despite this constraint, using our triplet-only contexts for gpt-4.1, it achieves strong results, reaching 64.5 EM and 76.0 F1 on HotpotQA and 46.8 EM and 59.8 F1 on MuSiQue. These scores surpass several retrieval-based methods, including ReadAgent and GraphReader, and are comparable to more resource-intensive systems such as AriGraph and HOLMES that rely on richer textual context. This demonstrates that complete and well-structured symbolic representations of KGs can serve as a sufficient and reliable information source for multi-hop reasoning.
@@ 4 Conclusions
We introduced Wikontic, a fully automated pipeline that uses LLMs to construct KGs from unstructured text. The pipeline produces compact, internally consistent graphs by aligning extracted triplets with the Wikidata ontology and deduplicating entities and relations. While using KGs as the sole knowledge source for multi-hop question answering, Wikontic achieves competitive performance with retrieval-augmented and KG-based baselines that still rely on source texts. Thus, ontology-guided KG construction is a viable alternative to passage-level retrieval. On MuSiQue, Wikontic includes 38–45 more unique entities than HippoRAG and AriGraph and contains the correct answer entity in 97.5% of cases. Within a 10-hop subgraph, it maintains 70% answer coverage and denser local connectivity. For QA, it attains 64.5 EM / 76.0 F1 on HotpotQA and 46.8 EM / 59.8 F1 on MuSiQue, outperforming text-reliant systems (ReadAgent, GraphReader) and approaching larger text-dependent methods (AriGraph, HOLMES). Moreover, Wikontic achieves state-of-the-art results on the MINE-1 benchmark, achieving 84–86% information-retention scores and surpassing GraphRAG and KGGen. These findings indicate that the proposed pipeline preserves a substantial amount of factual information from source texts.
@@
Our findings show that ontology-aware KG construction enables scalable, interpretable, and verifiable transformation of unstructured text into structured knowledge, bridging symbolic reasoning and generative language modeling.
@@
Our experiments are restricted to proprietary OpenAI models (GPT-4.1, GPT-4.1-mini, GPT-4o-mini) and the open-source Llama 3.3-70B. Token efficiency is measured as model-generated tokens during KG construction. This metric reflects provider billing but not end-to-end latency or throughput. Every stage of the pipeline currently uses LLMs with instructions and in-context examples prompting. Because the pipeline now yields its own annotated data, several stages could be replaced in future work by smaller, task-specific models fine-tuned on this data, thereby improving efficiency and lowering computational cost.
In the current work, we focus only on Wikidata due to its size, quality, and rich ontology. However, the proposed pipeline is flexible and can be adapted to any domain and ontology with a matching format of triplet constraints and entity types.
@@ A.1 Reproducibility
We release [^6] https://github.com/screemix/Wikontic: (1) prompts and source code for each pipeline stage; (2) scripts to build Wikidata-derived ontology used for ontology-aware stages; (3) the KG-only multi-hop QA component.
@@ A.2 Dataset Statistics
The test splits of HotpotQA and MuSiQue consist of 1000 samples each. We provide the QA evaluation results for Wikontic and HippoRAG for the whole evaluation set, and for AriGraph, we report the openly available results for 200 test samples. To compare the statistics of KGs, we use 80 MuSiQue samples that are commonly available for all compared methods.
@@ A.4 Prompts for triplet extraction
Here we provide excerpts of prompts that were used in our KG construction pipeline. A.4 was used for candidate triplet extraction. Subsequently, A.4 was used to refine entity types for both subject and object entities. A.4 was used for choosing relations among those that can legally connect chosen entity types by Wikidata constraints. Finally, A.5 was used to refine surface forms of subject and object. All prompts, instructions, and in-context examples are available with the code.
@@ A.5 Prompts for question answering
Here, we provide excerpts of prompts that were used for KG grounded question answering. A.5 was used to extract entities relevant to the question. Then, with A.5 the LLM was instructed to choose relevant entities among entities similar to the extracted ones in the KG. A.5 was used to decompose the question on a single-hop subquestion conditioned on extracted entities or previously answered subquestions. A.5 was used to check if a question is answered by a sequence of subquestions and corresponding answers. All prompts, instructions, and in-context examples are available with the code.